Circular No. 8218

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2000fw–2000ga AND 2001ki–2001ku D. S. Madgwick, Lawrence Berkeley Laboratory (LBL); P. C. Hewett,

D. J. Mortlock, University of Cambridge; and L. Wang, LBL, report the discovery of eighteen type-Ia supernovae detected spectroscopically in the first data release of the Sloan Digital Sky Survey (SDSS). The supernovae spectra were found to be superimposed over those of otherwise normal galaxies that had been targeted in this survey (details submitted to Ap.J. Letters). None of these objects were found in SDSS images, so that in each case, the tabulated positions are those of the host-galaxy centers, each supernova being within 3′′ of these positions in order to be observed through the narrow-field fiber apertures employed. The type and magnitude of each supernova has been estimated from template matches to the extracted supernova component of the observed spectrum. W. M. Wood-Vasey, LBL, reports that NEAT images (cf. IAUC 8150) within three months of the discovery dates given below were found for four fields, though no new object was visible to the following limiting magnitudes: 2001ki, 2001 Mar. 21, [20.5; 2001kk, Apr. 1, [19.5; 2001km, Apr. 7, [18.6; 2001kp, Mar. 16, [18.9.

SN 2000 UT α_{2000} δ_{2000} Mag. z
2000fw Mar. 10 14^{h}30^{m}14.^{s}07 + 0^{o}30′35′′.4 19.1  0.0962
2000fx Dec. 1 10 18 00.47 − 0 01 58.0 18.7 0.0652
2000fy Dec. 6 8 03 12.61 +47 36 49.7 19.5 0.1174
2000fz Dec. 15 1 18 35.83 +14 41 00.5 17.8 0.0539
2000ga Dec. 24 2 29 03.73 − 8 24 13.7 20.0 0.1408
SN 2001 UT α_{2000} δ_{2000} Mag. z
2001ki Jan. 1 10^{h}48^{m}58.^{s}40 − 0^{o}26′45′′.2 19.3 0.1062
2001kj Jan. 2 9 22 29.15 +57 54 29.3 18.1 0.0626
2001kk Jan. 19 12 47 33.41 + 0 05 57.1 18.8 0.0864
2001kl Jan. 22 3 08 51.44 − 1 10 24.1 20.4 0.1262
2001km Feb. 15 9 11 38.38 − 0 42 54.0 18.4 0.0702
2001kn Mar. 13 8 32 09.85 +47 17 27.6 19.9 0.1340
2001ko Mar. 16 14 10 58.32 +64 50 50.9 20.0 0.1406
2001kp Mar. 21 9 51 53.08 + 1 06 05.8 18.2 0.0634
2001kq Mar. 23 9 33 34.31 +55 10 26.0 19.1 0.0737
2001kr Mar. 26 9 59 15.75 + 0 58 02.4 19.1 0.0882
2001ks Apr. 18 7 56 46.50 +36 59 17.0 18.7 0.0780
2001kt May 25 16 17 13.40 +48 28 27.7 19.9 0.1038
2001ku June 19 13 39 44.77 +63 23 37.0 19.9 0.1358
2003 October 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green